OpenStudy (anonymous):

the sum of two numbers is 51. the smaller number is 23 less then the larger number. what are the numbers?

OpenStudy (anonymous):

Let the 2 numbers be x and y x+y=51 Let x be the smaller number and y be the larger number x=y-23 Now using substitution solve for y

OpenStudy (anonymous):

So x+y=51 We know that x=y-23 So lets replace x with y-23 (y-23)+y=51 Now lets solve for y y-23+y=51 2y-23=51 2y=74 y=37 Now we know that x=y-23 And we know that y=37 So lets replace y with 37 x=37-23 Solve for x BADDAAAMMM

OpenStudy (anonymous):

Thank you saved my life!!!